OpenStudy (anonymous):

The slope of a line is 5 and the point (16, -4) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

y = mx+b y=5x+b (as slope is 5) now plug in (16,-4) so -4=80 +b b=-84 hence y=5x-84

OpenStudy (amistre64):

y = 5(x-16)-4 y = 5x -80-4 y = 5x-84